A uniform plane wave has a wavelength of 6cm in vacuum and 2 cm in a non-magnetic dielectric having permeability of free space. Calculate the dielectric constant of the dielectric and the phase velocity of propagation of the wave through it.
1
Expert's answer
2019-05-23T09:39:55-0400

Wavelength in vacuum:


"\lambda_0 = \frac{c}{\nu} = 6\,\text{cm}"


Frequency "\nu" of both waves are equal.

In dielectric medium, speed of light equals to

"v = \frac{c}{n} = \frac{c}{\sqrt{\varepsilon \mu}}"

As dielectric is non-magnetic, wavelength equals to


"\lambda_m=\frac{v}{\nu} = \frac{c}{\nu\sqrt{\varepsilon}} = 2\,\text{cm}"

Therefore, dielectric constant


"\sqrt{\varepsilon} = \frac{\lambda_0}{\lambda_m} \Rightarrow \varepsilon = \frac{\lambda_0^2}{\lambda_m^2} = 9"

phase velocity of propagation of the wave:


"v = \frac{c}{\sqrt{\varepsilon }} = 10^8 \,\text{m/s}"